the length of a retangle is 3 less than 5 times of its width.
write a simplified algebraic expression for the perimeter of a rectangle.
if the rectangle width is tripled and its length is doubled the perimeter of new rectangle is 92 cm greater than original perimeter.
Find the area of the original rectangle.

Answer by mananth(16946)

let width =x
length = 5x-3
..
Perimeter = 2(x+5x-3)
Perimeter = 2(6x-3)
=12x-6
...
new rectangle
width = 3x
length = 2(5x-3)
perimeter = 2(3x+2(5x-3))
perimeter = 2(3x+10x-6)
perimeter = 6x+20x-12
...
6x+20x-12=12x-6+92
6x+20x-12x=-6+92+12
14x=98
/14
x=98/14
x=7 cm. the width of original rectangle
length = 5x-3 = 5*7-3 = 32cm
...
m.ananth@hotmail.ca